# Stability of Non-Hyperbolic Sets

### v1

@@ 4.2 The secondary tangencies
$dm_{6,y}$ $=$ $\left[D_{5}C\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+D_{5}Cn\left(\lambda^{\theta}\mu\right)^{2n}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]\right$.
$+$ $\left.O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right]dt$
$+$ $\left[D_{5}C\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+D_{5}C\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial a}\right$.
$+$ $\left.D_{5}Cn\left(\lambda^{\theta}\mu\right)^{2n}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]\right$.
$+$ $\left.O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right]da+O\left(\left(\lambda|^{\theta}\mu\right)^{{2n}}\right)dy,$
where, by (2.18)
$\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]>0.$
Refer to Figure 4. The formula for $m_{6,y}$ holds in general for all $\Delta t,\Delta a$ and $\Delta y$. However at the point $(t_{1},a)$ when $\Delta t=\Delta a=0$ we have ${\partial m_{6,y}}/{\partial y}=0$. Hence, the Taylor polynomial of second order for $\Delta m_{6,y}$ does not contain a linear term in $\Delta y$. As consequence ${\partial z^{(3)}_{y}}/{\partial t}={\partial m_{6,y}}/{\partial t}$ and ${\partial z^{(3)}_{y}}/{\partial a}={\partial m_{6,y}}/{\partial a}$. It is left to prove that $C,D_{5}$ converge. This is achieved since they are part of the derivative $DF^{N}$ which converges to $DF^{N}(0,1)$. The proposition follows. ∎
@@
Proposition 3.
For all $t\in\left(-t_{0},t_{0}\right)$ and for $n$ large enough there exists $t_{n}\in\left(-t_{0},t_{0}\right)$ such that the following holds:
    $|t-t_{n}|=O\left(\frac{1}{n}\right)$,
    $F_{t_{n},a_{n}(t_{n})}$ has a secondary tangency $q_{1,n,n_{0}}$.
@@
Choose $t^{*}\in\left(-t_{0},t_{0}\right)$. Use the notation of Proposition 2 and notice that, by the choice of $\theta$, see (2.18),
$DC^{2}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]=v>0$
@@
Take $n$ large enough and a point $(t,a_{n}(t))\in A_{n}$ near $(t^{*},a_{n}(t^{*}))\in A_{n}$. By Proposition 2 and Lemma 3
$z^{(3)}_{y}\left(t,a_{n}(t)\right)$ $=$ $z^{(3)}_{y}\left(t^{*},a_{n}(t^{*})\right)+\int_{0}^{(t-t^{*})}\left[\frac{\partial z^{(3)}}{\partial t}+\frac{\partial z^{(3)}}{\partial a}\frac{da_{n}}{dt}\right]dt=z^{(3)}_{y}\left(t^{*},a_{n}(t^{*})\right)$ (4.48)
$+$ $\left\{DC\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+DC^{2}n\left(\lambda^{\theta}\mu\right)^{2n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]\right$.
- $\left.DC\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right\}(t-t^{*})$
$=$ $z^{(3)}_{y}\left(t^{*},a_{n}(t^{*})\right)+n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(t-t^{*}),$
@@
where we used that $\lambda^{2\theta}\mu^{3}>1$ and $\lambda^{\theta}\mu^{2}>1$ see (2.16). Observe that we have a cancellation of the dominant terms in the partial derivatives obtained by combining Proposition 2 and Lemma 3. Let $n^{*}_{0}$ be such that $W^{u}_{\text{loc}}\left(z^{(3)}\left(t^{*},a_{n}(t^{*})\right)\right)\subset S_{n-n_{0}^{*}}$, then, by (2.8),
$z^{(3)}_{y}\left(t^{*},a_{n}(t^{*})\right)\geq\frac{2}{\mu}\frac{1}{\mu^{n-n^{*}_{0}}}.$ (4.49)
By (4.48) and (4.49)
$z^{(3)}_{y}\left(t,a_{n}(t)\right)\geq\frac{2}{\mu}\frac{1}{\mu^{n-n^{*}_{0}}}+n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(t-t^{*}).$ (4.50)
@@
Choose $\kappa\geq 2$. Then by (2.8), in a neighborhood of $(z^{(3)}\left(t^{*},a_{n}(t^{*})\right),t^{*},a_{n}(t^{*}))$
$\max W_{n-n_{0}^{*}-\kappa}$ $\leq$ $\frac{2}{\mu^{n-n^{*}_{0}-\kappa}}.$
From the previous inequality, (4.50), we get that $W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))$ is above $W_{n-n_{0}^{*}-\kappa}$ if
$n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(t-t^{*})\geq\frac{2}{\mu^{n-n^{*}_{0}}}\left[\mu^{\kappa}-\frac{2}{\mu}\right]\geq 2C^{\prime}\left(\lambda^{\theta}\mu\right)^{2n}\left[\mu^{\kappa}-\frac{2}{\mu}\right],$
where we also used (4.17) and $C^{\prime}$ is a uniform constant. As consequence, if
$t-t^{*}\geq\frac{4C^{\prime}}{n}\frac{\mu^{\kappa}}{v},$
@@
then, $W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))$ is above $W_{n-n_{0}^{*}-\kappa}$. Because $W^{u}_{\text{loc}}(z^{(3)}(t^{*},a_{n}(t^{*})))$ contains a point below $W_{n-n_{0}^{*}-1}$, there exists a parameter between $(t,a_{n}(t))$ and $(t^{*},a_{n}(t^{*}))$ for which a secondary homoclinic tangency of type $n_{0}=n_{0}^{*}+\kappa$ occurs and $|t^{*}-t|=O\left(\frac{1}{n}\right)$. Moreover by Lemma 2, (2.8), Lemma 11 and the definition of $\alpha$ we get
$\frac{\partial W_{n-n_{0}}}{\partial t}(q_{1,n,n_{0}})=-2n\left(\lambda^{\theta}\mu\right)^{2n}\mu^{\kappa-1}\left(1-\alpha-\frac{\kappa}{n}\right)\frac{\partial\mu}{\partial t}+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right).$ (4.51)
Observe that, from $(F6)$, the value of $\alpha$ and the lower bound for $\theta$, see (2.17), we have
$\left(1-\alpha-\frac{\kappa}{n}\right)\frac{\partial\mu}{\partial t}>0.$ (4.52)
@@
Use coordinates centered at $z^{(2)}$ and let $(\Delta x,\Delta y)\in W^{u}_{\text{loc}}(z^{(2)})$ such that $F^{N}_{t,a}(\Delta x,\Delta y)=q_{1,n,n_{0}}$. Then $0\leq\Delta y\leq L\left(\lambda^{\theta}\mu\right)^{n}$. Let $v=\left(\begin{matrix}v_{1}\\ v_{2}\end{matrix}\right)$ be the tangent vector at $(\Delta x,\Delta y)$ to $W^{u}_{\text{loc}}(z^{(2)})$. We use Lemma 8 and we obtain that
$|v_{2}|=\left[2Q_{2}\left(\frac{\mu}{\lambda^{2}}\right)^{n}|\Delta x|+O\left(|\Delta x|^{2}\left(\frac{\mu}{\lambda^{3}}\right)^{n}\right)\right]|v_{1}|.$ (4.53)
By Lemma 1 we have that ${\partial^{2}Y}/{\partial y^{2}}$ is bounded away from zero in a neighborhood of $c$. As consequence, by Lemma 7, Lemma 7 and Lemma 5 we have $DF^{N}_{t,a}(\Delta x,\Delta y)=\left(\begin{matrix}A&B\\ C&D\end{matrix}\right)$ where $D\geq d\left(\lambda^{\theta}\mu\right)^{n}$ and $d>0$ is a uniform constant. Because $F^{N}_{t,a}(\Delta x,\Delta y)$ is a tangency at $q_{1,n,n_{0}}$, then
$Cv_{1}+Dv_{2}=O\left(\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\left(|v_{1}|+|v_{2}|\right)\right),$ (4.54)
where we used (2.9). The proof of the lemma is completed by following exactly that of Lemma 10. ∎
@@ 5 Curves of secondary tangencies
In the previous section, we proved the existence of tangency points on the curve $A_{n}$. In the sequel, we extend these points to create global curves. All points along these curves have a secondary tangency; refer to Figure 6. In particular, we prove that the curves of secondary tangencies are actually long uniform curves that never leave the strip $\mathcal{B}_{n}$.
@@
5.1 Continuation of secondary tangencies: the implicit function theorem
The next lemma is the first step in proving the existence of local tangency curves. In particular, we prove that each parameter in $\mathcal{B}_{n}$ corresponding to a map having a secondary tangency of type $n_{0}$ can be locally enlarged to create a curve. Each point on this curve corresponds to a map with a $n_{0}$ secondary tangency. This result is achieved by implementing and using the implicit function theorem.
Lemma 13.
Let $(t^{*},a^{*})\in{\mathcal{B}_{n}}$ be a parameter corresponding to a map with a secondary tangency of type $n_{0}$. Then there exist $\epsilon>0$ and a smooth curve
$b_{n,n_{0}}:[t^{*}-\epsilon,t^{*}+\epsilon]\mapsto\mathbb{R}$
such that, for all $t\in[t^{*}-\epsilon,t^{*}+\epsilon]$ the parameter $(t,b_{n,n_{0}}(t))$ corresponds to a map with a secondary tangency of type $n_{0}$. Moreover
$\frac{(C-E)}{T}\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}+O\left(n\lambda^{\theta n}\right)\leq\frac{db}{dt}\leq-\frac{T}{(C-E)}\frac{n}{\mu^{n+1}}\frac{\partial\mu}{\partial t}+O\left(\frac{n}{\mu^{n_{0}/2}}\right).$
In particular, there exists a positive constant $v$ such that,
$\frac{db}{dt}=\frac{da_{n}}{dt}+v{n\lambda^{\theta n}}+O\left(\lambda^{\theta n}\right)$
for all $t\in[\tilde{t}^{*}-\epsilon,\tilde{t}^{*}+\epsilon]$, with $(\tilde{t}^{*},b_{n,n_{0}}(\tilde{t}^{*}))\in a_{n}$.
@@ 5.1 Continuation of secondary tangencies: the implicit function theorem
Let $(t^{*},a^{*})\in\mathcal{B}_{n}\cap A_{n}$ be a parameter corresponding to a map with a secondary tangency of type $n_{0}$, $q_{1,n,n_{0}}$. Its existence is proved in Proposition 3. We construct a local function whose graph is contained in a neighborhood of $t^{*}$. Let $m_{1}\in W^{u}_{\text{loc}}(c)$ such that $F_{t,a}^{3N+(1+\theta)n}(m_{1})=q_{1,n,n_{0}}$. To describe the perturbation of $W^{u}_{\text{loc}}(q_{1,n,n_{0}})$ we define the following function by choosing coordinates centered in the image at $q_{1,n,n_{0}}$. Take some $\epsilon>0$ and consider the ${\mathcal{C}}^{4}$} function $(\tilde{x},\tilde{y}):(-\epsilon,\epsilon)^{3}\mapsto\mathbb{R}^{m-1}\times\mathbb{R}$ defined by
$\left(\tilde{x}(\Delta y,\Delta t,\Delta a),\tilde{y}(\Delta y,\Delta t,\Delta a)\right)=F_{t+\Delta t,a+\Delta a}^{3N+(1+\theta)n}\left(m_{1}+(0,\Delta y)\right),$
which describes $W^{u}_{\text{loc}}\left(F_{t+\Delta t,a+\Delta a}^{3N+(1+\theta)n}(m_{1})\right)$. Observe that the manifolds $W_{n-n_{0}}$ of $F_{t+\Delta t,a+\Delta a}$ are described locally, near $q_{1,n,n_{0}}$, as the graph of a ${\mathcal{C}}^{4}$} function

### v3

@@ 4.2 The secondary tangencies
$dm_{6,y}$ $=$ $\left[D_{5}S_{C}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+D_{5}S_{C}Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]\right$.
$+$ $\left.O\left(S_{C}\left(\lambda^{\theta}\mu\right)^{n}\right)\right]dt$
$+$ $\left[D_{5}S_{C}\mu^{n}+1+D_{5}S_{C}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+D_{5}S_{C}Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{1}{\mu}\frac{\partial\mu}{\partial a}\right]\right$.
$+$ $\left.O\left(S_{C}\left(\lambda^{\theta}\mu\right)^{n}\right)\right]da+O\left(\left(\lambda^{\theta}\mu\right)^{{2n}}\right)dy$ (4.2.31)
where $S_{C}=C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left(\lambda^{\theta}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)$, and by (2.3.3)
$\left|\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right|>w>0.$
The formula for $m_{6,y}$ holds in general for all $\Delta t,\Delta a$ and $\Delta y$. However at the point $(t_{1},a)$ when $\Delta t=\Delta a=0$ we have ${\partial m_{6,y}}/{\partial y}=0$. Hence, the Taylor polynomial of second order for $\Delta m_{6,y}$ does not contain a linear term in $\Delta y$. As consequence ${\partial z^{(3)}_{y}}/{\partial t}={\partial m_{6,y}}/{\partial t}$ and ${\partial z^{(3)}_{y}}/{\partial a}={\partial m_{6,y}}/{\partial a}$. It is left to prove that $C,D_{5}$ converge. This is achieved since they are part of the derivative $DF^{N}$ which converges to $DF^{N}(0,1)$. The proposition follows. ∎
@@
Proposition 4.2.32.
For all $t\in\left(-t_{0},t_{0}\right)$ and for $n$ large enough there exists $t_{n}\in\left(-t_{0},t_{0}\right)$ such $t-t_{n}=O\left({1}/{n}\right)$ and $F_{t_{n},a_{n}(t_{n})}$ has a secondary tangency.
@@
Choose $t\in\left(-t_{0},t_{0}\right)$. Use the notation of Proposition 4.2.2 and notice that, by the choice of $\theta$, see (2.3.3),
$\left|\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right|>w>0.$
Suppose that,
$DC^{2}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]=v>0,$
@@
the negative case is handled similarly by reversing the time. Take $n$ large enough and a point $(s,a_{n}(s))\in A_{n}$ near $(t,a_{n}(t))\in A_{n}$. By Proposition 4.2.2 and Lemma 2.2.6
$z^{(3)}_{y}\left(s,a_{n}(s)\right)$ $=$ $z^{(3)}_{y}\left(t,a_{n}(t)\right)+\int_{0}^{(s-t)}\left[\frac{\partial z^{(3)}}{\partial t}+\frac{\partial z^{(3)}}{\partial a}\frac{da_{n}}{dt}\right]dt=z^{(3)}_{y}\left(t,a_{n}(t)\right)$ (4.2.33)
$+$ $DS_{C}\left\{\frac{n}{\mu}\frac{\partial\mu}{\partial t}+Cn\left(\lambda^{\theta}\mu\right)^{n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]-\frac{n}{\mu}\frac{\partial\mu}{\partial t}+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right\}(s-t)$
$=$ $z^{(3)}_{y}\left(t,a_{n}(t)\right)+\left\{DC^{2}n\left(\lambda^{\theta}\mu\right)^{2n}\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right\}(s-t)$
$\geq$ $z^{(3)}_{y}\left(t,a_{n}(t)\right)+n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(s-t),$
@@
where we used that $\lambda^{2\theta}\mu^{3}>1$ and $\lambda^{\theta}\mu^{2}>1$ see (2.3.1). Observe that we have a cancellation of the dominant terms in the partial derivatives obtained by combining Proposition 4.2.2 and Lemma 2.2.6. Let $n_{0}=n_{0}(t,a_{n}(t))$ be as in (4.1.46). Hence, by (2.2.2),
$z^{(3)}_{y}\left(t,a_{n}(t)\right)\geq\frac{2}{\mu}\frac{1}{\mu^{n-n_{0}}}.$ (4.2.34)
By (4.2.33) and (4.2.34)
$z^{(3)}_{y}\left(s,a_{n}(s)\right)\geq\frac{2}{\mu}\frac{1}{\mu^{n-n_{0}}}+n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(s-t).$ (4.2.35)
@@
Choose $\kappa\geq 2$ such that $\left[\mu^{\kappa}-{2}/{\mu}\right]\geq v/4$. Then by (2.2.2), in a neighborhood of $(z^{(3)}\left(t,a_{n}(t)\right),t,a_{n}(t))$
$\max W_{n-n_{0}-\kappa}$ $\leq$ $\frac{2}{\mu^{n-n_{0}-\kappa}}.$
From the previous inequality, (4.2.35), we get that $W^{u}_{\text{loc}}(z^{(3)}(s,a_{n}(s)))$ is above $W_{n-n_{0}^{*}-\kappa}$ if
$n\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2}(s-t)\geq\frac{2}{\mu^{n-n_{0}}}\left[\mu^{\kappa}-\frac{2}{\mu}\right]\geq 2C^{\prime}\left(\lambda^{\theta}\mu\right)^{2n}\left[\mu^{\kappa}-\frac{2}{\mu}\right]\geq C^{\prime}\left(\lambda^{\theta}\mu\right)^{2n}\frac{v}{2},$
where we also used (4.1.49) and $C^{\prime}$ is a uniform constant. As consequence, if
$s-t\geq\frac{C^{\prime}}{n},$
@@
then, $W^{u}_{\text{loc}}(z^{(3)}(s,a_{n}(s)))$ is above $W_{n-n_{0}-\kappa}$. Because $W^{u}_{\text{loc}}(z^{(3)}(t,a_{n}(t)))$ contains a point below $W_{n-n_{0}-1}$, there exists a parameter between $(s,a_{n}(s))$ and $(t,a_{n}(t))$ for which a secondary homoclinic tangency of type $n_{0}+1$ occurs and $t-s=O\left({1}/{n}\right)$. ∎
@@
Use coordinates centered at $z^{(3)}$ and let $(\Delta x_{z^{(3)}},\Delta y_{z^{(3)}})\in W^{u}_{\text{loc}}(z^{(3)})$ corresponding to the tangency point $q_{1,n,n_{0}}$. Then by (2.2.3),
$\frac{d\Delta y_{z^{(3)}}}{d\Delta x_{z^{(3)}}}=O\left(\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\right)$
and by differentiating the equation in Lemma 4.1.34,
$\frac{d\Delta y_{z^{(3)}}}{d\Delta x_{z^{(3)}}}=O\left(\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\right)=\frac{16Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\Delta x_{z^{(3)}}+O\left(\Delta x_{z^{(3)}}^{2}\right).$ (4.2.37)
In particular,
${\Delta x_{z^{(3)}}}=O\left(\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\left(\frac{\lambda^{2}}{\mu}\right)^{n}\frac{1}{H^{3}}\right).$
The curvature at the tangency point is found by differentiating (4.2.37). Hence,
$\frac{d^{2}\Delta y_{z^{(3)}}}{d\Delta x_{z^{(3)}}^{2}}$ $=$ $\frac{16Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}+O\left(\Delta x_{z^{(3)}}\right)$
$\geq$ $\frac{16Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}\left[1-O\left(\left(\frac{\lambda}{\mu}\right)^{n-n_{0}}\left(\frac{\lambda^{4}}{\mu^{2}}\right)^{n}\frac{1}{H^{6}}\right)\right]$
$\geq$ $\frac{8Q^{4}}{C^{2}B^{4}}{H^{3}}\left(\frac{\mu}{\lambda^{2}}\right)^{n}.$
where we used the expression of $H$ given in Lemma 4.1.23 and the lower bound of the strip $\mathcal{B}_{n}$. The lower bound for $H$ gives the statement of the Lemma. ∎
@@ 5 Curves of secondary tangencies
In the previous section, we proved the existence of maps with secondary tangency points on the curve $a_{n}$. In the sequel, we show that secondary tangencies persists along curves of uniform size. In particular these curves are graphs of functions defined on all of the $t$ domain, see Figure 6. Moreover, we prove that the curves of secondary tangencies are actually long uniform curves that are actually contained in the strip $\mathcal{B}_{n}$.
@@
5.1 Persistence of secondary tangencies
The next lemma is the first step in proving the existence of local tangency curves. In particular, we prove that each parameter in $\mathcal{B}_{n}$ corresponding to a map having a secondary tangency of type $n_{0}$ is contained in a local curve. Each point on this curve corresponds to a map with a $n_{0}$ secondary tangency.
Lemma 5.1.1.
Let $(t_{n_{0}},a_{n_{0}})\in{\mathcal{B}_{n}}$ be a parameter corresponding to a map with a secondary tangency of type $n_{0}$. Then there exist $\epsilon>0$ and a smooth curve
$b_{n,n_{0}}:[t_{n_{0}}-\epsilon,t_{n_{0}}+\epsilon]\mapsto\mathbb{R}$
such that, for all $t\in[t_{n_{0}}-\epsilon,t_{n_{0}}+\epsilon]$ the parameter $(t,b_{n,n_{0}}(t))$ corresponds to a map with a secondary tangency of type $n_{0}$. Moreover, the slopes of the curves $b_{n,n_{0}}$ tend exponentially fast to zero, namely
$\left|\frac{db_{n,n_{0}}}{dt}\right|=O\left(\frac{n}{\left(\lambda^{\theta}\mu^{2}\right)^{n}\mu^{n_{0}^{*}/2}}\right).$
In particular,
$\frac{db_{n,n_{0}}}{dt}=\frac{da_{n}}{dt}+{n\lambda^{\theta n}}C\left[\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{1}{\mu}\frac{\partial\mu}{\partial t}\right]+O\left(\lambda^{\theta n}\right),$
for all $t\in[t_{n_{0}}-\epsilon,t_{n_{0}}+\epsilon]$, with $(t,b_{n,n_{0}}(t))\in a_{n}$.
@@ 5.1 Persistence of secondary tangencies
First we consider the case when the given point $(t_{n_{0}},a_{n_{0}})\in\mathcal{B}_{n}$ is on the curve $a_{n}$, i.e. $a_{n_{0}}=a_{n}(t_{n_{0}})$. In particular this point corresponds to a map with a secondary tangency of type $n_{0}$, $q_{1,n,n_{0}}$. We construct a local function whose graph contains the point $(t_{n_{0}},a_{n_{0}})$. Let $m_{1}\in W^{u}_{\text{loc}}(c)$ such that $F_{t,a}^{3N+(1+\theta)n}(m_{1})=q_{1,n,n_{0}}$. To describe the perturbation of $W^{u}_{\text{loc}}(q_{1,n,n_{0}})$ we define the following function by choosing coordinates centered in $q_{1,n,n_{0}}$. Take some $\epsilon>0$ and consider the ${\mathcal{C}}^{4}$} function $(\tilde{x},\tilde{y}):(-\epsilon,\epsilon)^{3}\mapsto\mathbb{R}\times\mathbb{R}$ defined by
$\left(\tilde{x}(\Delta y,\Delta t,\Delta a),\tilde{y}(\Delta y,\Delta t,\Delta a)\right)=F_{t_{n_{0}}+\Delta t,a_{n_{0}}+\Delta a}^{3N+(1+\theta)n}\left(m_{1}+(0,\Delta y)\right),$
which describes $W^{u}_{\text{loc}}\left(F_{t_{n_{0}}+\Delta t,a_{n_{0}}+\Delta a}^{3N+(1+\theta)n}(m_{1})\right)$. Observe that the manifolds $W_{n-n_{0}}$ of $F_{t_{n_{0}}+\Delta t,a_{n_{0}}+\Delta a}$ are described locally, near $q_{1,n,n_{0}}$, as the graph of a ${\mathcal{C}}^{4}$} function